Close to 15 hospitals in Maharashtra has to be completely or partially shut down after
many of their staff were quarantined following exposure to COVID-19 patients. The
number x of medical staff entering the hospitals treating COVID-19 patients on any
week has a Poisson probability distribution with average equal to 5 staff per week.
(a) What is the probability that the number of staff entering the hospital on a particular
week is two
(b) Is it likely that x will exceed 10? Substantiate your answer.
1
Expert's answer
2020-08-03T19:13:30-0400

(a) Poisson probability


"P(X=x)=\dfrac{e^{-\mu}\cdot \mu^x}{x!}"

Given "\mu=5"


"P(X=2)=\dfrac{e^{-5}\cdot 5^2}{2!}\approx0.084224"

(b)


"P(X>10)=1-P(X\leq 10)=""=1-P(X=0)-P(X=1)-P(X=2)-""-P(X=3)-P(X=4)-P(X=5)-""-P(X=6)-P(X=7)-P(X=8)-""-P(X=9)-P(X=10)\approx0.0137"

We note "P(X>10)=0.0137<0.05." We consume that the probability "P(X>10)=0.0137" is small, which indicates that the event is unlikely to occur by chance and thus it is not likely that "X" will exceed 10.